### v1

# Deciphering interventional dynamical causality from non-intervention systems

[…]
-3pt plus 3pt minus 2pt
Deciphering interventional dynamical causality from non-intervention systems
Jifan Shi1,2,3, Yang Li4, Juan Zhao5, Siyang Leng1,6, Kazuyuki Aihara4∗, Luonan Chen4,7,8∗, Wei Lin1,2,3∗
1Research Institute of Intelligent Complex Systems, Fudan University, Shanghai 200433, China
2Shanghai Artificial Intelligence Laboratory, Shanghai 200232, China
3State Key Laboratory of Medical Neurobiology and MOE Frontiers Center for Brain Science, Fudan University, Shanghai 200032, China
4International Research Center for Neurointelligence, The University of Tokyo Institutes for Advanced Study, The University of Tokyo,Tokyo 113-0033, Japan
5School of Pharmacy, Shanghai University of Traditional Chinese Medicine, Shanghai 201203, China
6Academy for Engineering and Technology, Fudan University, Shanghai 200433, China
7Key Laboratory of Systems Biology, Shanghai Institute of Biochemistry and Cell Biology, Center for Excellence in Molecular Cell Science, Chinese Academy of Sciences, Shanghai 200031, China
8Key Laboratory of Systems Health Science of Zhejiang Province, School of Life Science, Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Chinese Academy of Sciences, Hangzhou 310024, China
To whom correspondence should be addressed; E-mail: kaihara@g.ecc.u-tokyo.ac.jp or lnchen@sibs.ac.cn or wlin@fudan.edu.cn.
Detecting and quantifying causality is a focal topic in the fields of science, engineering, and interdisciplinary studies. However, causal studies on non-intervention systems attract much attention but remain extremely challenging. To address this challenge, we propose a framework named Interventional Dynamical Causality (IntDC) for such non-intervention systems, along with its computational criterion, Interventional Embedding Entropy (IEE), to quantify causality. The IEE criterion theoretically and numerically enables the deciphering of IntDC solely from observational (non-interventional) time-series data, without requiring any knowledge of dynamical models or real interventions in the considered system. Demonstrations of performance showed the accuracy and robustness of IEE on benchmark simulated systems as well as real-world systems, including the neural connectomes of C. elegans, COVID-19 transmission networks in Japan, and regulatory networks surrounding key circadian genes.
Introduction
Directional or indirectional interactions among different components give rise to a variety of complex phenomena in nature and social society. The causality or the causal effect is one of the most attractive directional relations originating from the non-reversibility of time. Modeling, detecting, and quantifying causality from observational data are crucial for describing, interpreting, predicting, and even controlling complex systems.
Statisticians believe that the causality is contained and obtainable in universal random variables regardless of the time label. The Neyman-Rubin’s potential outcome framework  (?, ?, ?), the Wright’s structural equation model  (?, ?) and the Pearl’s causal diagram model  (?, ?) are the most famous statistical approaches, which have been proved to be of mathematical equivalence  (?). Especially, the instrumental variable method, that was awarded the 2021 Nobel Prize in economic sciences, is widely applied in the study of causality among economics, environments, and other disciplines  (?). Statistical methods seek to discern binary causal relations among random variables on a directed acyclic graph, without relying on temporal data. However, in general complex dynamical systems, the causation must precede the effect, feedback is common, and the causal strength needs to be quantitatively measured. Time should play definitely a key role.
Dynamics-oriented researchers have proposed fruitful algorithms for measuring the observational causality from time series, commonly referred to as the dynamical causality  (?). The celebrated Granger causality (GC)  (?, ?) employs a linear model and uses improvement of predictability over time to illustrate the causality. Transfer entropy (TE)  (?) generalizes GC to the nonlinear case by quantifying the prediction uncertainty through Shannon entropy. Neither GC nor TE addresses the “non-separability” problem, which means that removing the causal variable from the system inevitably influences the dynamics of downstream variables  (?, ?, ?, ?). To measure causality in non-separable systems, numerous approaches have emerged in the last decade within the framework of delay embedding. These include convergence of cross mapping (CCM)  (?, ?), partial cross mapping (PCM)  (?), continuity scaling  (?), inverse continuity  (?), topological expansion  (?), joint distance distribution  (?), embedding entropy (EE)  (?) and other indices  (?, ?).
Nonetheless, most indices designed for quantifying the dynamical causality primarily estimate directional causal relations at the observational level, i.e. observational dynamical causality (ObsDC). Conventionally, to detect and quantify causality at the interventional level, which is noted as the interventional dynamical causality (IntDC) in this study, intelligent modulation and manipulation of the dynamical system should be taken into account. By allowing external intervention to the system and recording data under different perturbations, frameworks such as perturbation cascade inference (PCI)  (?) and dynamical causal effect (DCE)  (?, ?) provide relevant computational schemes. However, due to ethical or practical limitations, many real systems should be analyzed without any external intervention. An essential problem is how to measure IntDC solely from the observational data but at the interventional level.
In this study, we propose the IntDC framework and introduce a criterion named Interventional Embedding Entropy (IEE), which aims to identify and quantify IntDC between variables solely from the observational time series. Actually, IEE is rigorously derived by the theory of the delay embedding. Numerical IEE does not require specific prior knowledge of dynamics, and additional perturbation to the system is also unnecessary. Compared to ObsDC indices, IEE designed for IntDC has the capability to rank the importance of causal effects and construct directional causal networks more effectively. We demonstrate numerical experiments on both simulated examples and real datasets, including estimating neural connectomes of C. elegans, evaluating COVID-19 transmission in Japan, and constructing regulatory networks surrounding key circadian genes.
Methods
Observational dynamical causality in delay-embedding space
We consider a discrete two-variable dynamical system as
$\begin{cases}x_{t+1}=g(x_{t},x_{t-1},\dots,x_{t-p},\varepsilon_{x,t}),\\ y_{t+1}=f(x_{t},x_{t-1},\dots,x_{t-p},y_{t},y_{t-1},\dots,y_{t-p},\varepsilon_{y,t}),\end{cases}$ (1)
where $x,y$ are two variables, $p$ denotes the time step during which causality is considered, and $\varepsilon_{\cdot,t}$ stand for small noise terms. According to Shi et al.  (?), there exists ObsDC from $x$ to $y$ since the evolving equation of $y_{t+1}$ depends on the historical behavior of $x$, while there is no ObsDC from $y$ to $x$ as the dynamics of $x_{t+1}$ is independent of $y$. GC and TE measure the predictability of $f$ in the original time-series space to quantify the ObsDC from $x$ to $y$ through linear regression and entropy uncertainty, respectively (Fig. 1(B)). The causal strength of GC/TE is based on the assumption of separability, which means that removing $x$ from the system does not influence the observed data of $y$. However, in general coupled systems, the separability is not satisfied, as removing one variable changes the values of affected ones  (?, ?, ?).
To address the challenge of detecting ObsDC in universal non-separable systems, various approaches from the delay-embedding space have been proposed. In the delay-embedding framework, the bivariate system Eq. (1) is assumed to evolve into an attractive manifold with an inner dimension $d$. Let the time-delay vectors of $x$ and $y$ be
$\bm{X}_{t}=(x_{t},x_{t-1},\dots,x_{t-L})^{T}\in\mathcal{M}_{X}\subseteq\mathbb{R}^{L+1},$ (2)
and
$\bm{Y}_{t+1}=(y_{t+1},y_{t},y_{t-1},\dots,y_{t-L+1})^{T}\in\mathcal{M}_{Y}\subseteq\mathbb{R}^{L+1},$ (3)
where $L$ is the time-delay length, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ represent the manifolds formed by $\bm{X}_{t}$ and $\bm{Y}_{t}$, respectively. According to the seminal stochastic version of Takens’ embedding theorem  (?, ?, ?, ?), for delay dimension $L\geqslant 2d$, we can obtain
$\bm{X}_{t}=\bm{F}(\bm{Y}_{t+1}),$ (4)
where $\bm{F}$ is a smooth projection operator. The detailed derivation can be referred to the Supplementary Text. Equation (4) shows that the causal variable $\bm{X}_{t}\in\mathcal{M}_{X}$ can be reconstructed by the effect variable $\bm{Y}_{t+1}\in\mathcal{M}_{Y}$. However, as $\bm{F}$ is not reversible in generic sense, we can not determine $\bm{Y}_{t+1}$ only by the information from $\bm{X}_{t}$. The causal dependence in $f$ between $x$ and $y$ is transformed into the reconstructability of $\bm{F}$ (Fig. 1(D)). Instead of fitting a model by removing the causal variable as in GC and TE, detecting the existence and quantifying the continuity characteristics of $\bm{F}$ in the delay-embedding space can be adequate for causal identification, especially for universal non-separable systems. Related algorithms include CCM, PCM, and EE.
Figures 1(A), (B) and (D) summarize dynamical causality at the observational level, where the ObsDC from variable $x$ to $y$ is estimated only from the historical observed time series. But fitting Eq. (1) or Eq. (4) from data is criticized as mere association or prediction  (?, ?).
Figure 1: Illustration for the observational dynamical causality (ObsDC) and interventional dynamical causality (IntDC). (A) An illustrative example of a complex system where $(x_{t},y_{t})$ are observed time series. The $z_{t}$ stands for the unobserved variable. (B) ObsDC from $x$ to $y$ indicates that the dynamics of $y$ depends on $x$. Algorithms such as Granger causality (GC) and transfer entropy (TE) infer the ObsDC in the original time-series space. (C) IntDC from $x$ to $y$ indicates that the deviation of $y$, i.e. $\delta y$, depends on an intervention on $x$, i.e. $\delta x$. Usually additional data from a perturbed system $(\tilde{x}_{t},\tilde{y}_{t})$ is necessary for detecting IntDC. (D) ObsDC can be detected in the embedding space by cross mapping $\mathbb{F}$. The causal vector $X_{t}$ can be reconstructed by the effect vector $\bm{Y}_{t+1}$. (E) IntDC can be modeled in the embedding space. The interventional causal vector $\delta\bm{X}_{t}$ should be reconstructable by the effect vector $\delta\bm{Y}_{t+1}$ around $\bm{Y}_{t+1}$. The $\delta\bm{Y}_{t+1}$ is approximated by the distance from $\bm{Y}_{t+1}$ to its neighbors $\bm{Y}_{t_{k}+1}$, while $\delta\bm{X}_{t}$ is obtained by $X_{t_{k}}$ with the same time label $t_{k}$. (F) Interventional embedding entropy (IEE) measures the average information retention contained in $\delta\bm{Y}_{t+1}$ from $\delta\bm{X}_{t}$, quantifying the IntDC from variable $x$ to $y$. Notably, IEE does not require additional perturbations to the system and can be inferred from observational data solely.
Interventional dynamical causality in delay-embedding space
The detection of hidden essential-level causality necessitates intervention or manipulation to the system  (?, ?, ?, ?). We define that there is an IntDC from $x$ to $y$ during the time $[t-p,t]$ in the original time-series space, if $\delta y_{t}=\tilde{y}_{t}-y_{t}$ depends on $\delta x_{t}=\tilde{x}_{t}-x_{t}$, where the dynamics of $x_{t}$ is perturbed to $\tilde{x}_{t}$, and $\tilde{y}_{t}$ is the dynamics of affected $y$ after the perturbation (Fig. 1(C)). The intervention $\delta x_{t}$ can be either a pulse stimulation or persistent disturbances to the system.
To measure the IntDC, data before and after the intervention, i.e. $(x_{t},y_{t})$ and $(\tilde{x}_{t},\tilde{y}_{t})$, are typically required, as in PCI and DCE. These conventional algorithms rely on reproducible dynamics or data availability under different settings.
To address challenges of causal inference in widespread non-intervention, non-linear and non-separable systems, we turn to the study of intervened dynamics Eq. (4) in the delay-embedding space. Denote $\delta\bm{X}_{t}=\widetilde{\bm{X}}_{t}-\bm{X}_{t}$ and $\delta\bm{Y}_{t+1}=\widetilde{\bm{Y}}_{t+1}-\bm{Y}_{t+1}$, where $\widetilde{\bm{X}}_{t}$ and $\widetilde{\bm{Y}}_{t+1}$ are the time-delay vectors of $x$ and $y$ after the intervention, respectively. Under infinitesimal intervention, Eq. (4) leads to
$\delta\bm{X}_{t}=\nabla_{\bm{Y}}\bm{F}(\bm{Y}_{t+1})\cdot\delta\bm{Y}_{t+1}.$ (5)
Equation (5) indicates that $\delta\bm{Y}_{t+1}$ contains complete information of $\delta\bm{X}_{t}$ in the neighborhood of $\bm{Y}_{t+1}$ or reconstructs $\delta\bm{X}_{t}$, if there exists IntDC from the causal variable $x$ to the effect variable $y$ (Fig. 1(E)).
Figures 1(A), (C), and (E) give an overview of the IntDC framework.
Interventional embedding entropy
We introduce the IEE criterion, specifically designed for quantifying IntDC solely from the observational data in the delay-embedding space. The IEE criterion is proposed to measure the average information retention or reconstructability from the effect $\delta\bm{Y}_{t+1}$ to the causation $\delta\bm{X}_{t}$, formulated as
$\text{IEE}[x\rightarrow y]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}),$ (6)
where CMI is the conditional mutual information  (?), i.e.
$\text{CMI}(\bm{x},\bm{y}|\bm{z})=\iiint p(\bm{x},\bm{y},\bm{z})\log\frac{p(\bm{x},\bm{y}|\bm{z})}{p(\bm{x}|\bm{z})p(\bm{y}|\bm{z})}\mathop{}\!\mathrm{d}\bm{x}\mathop{}\!\mathrm{d}\bm{y}\mathop{}\!\mathrm{d}\bm{z}.$ (7)
The $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs. 1(E) and (F)). Thus, Eq. (6) can be numerically approximated only from the observational data, especially for non-intervention systems. Table 1 shows the numerical algorithm for IEE, whose detailed derivation can be found in the Supplementary Text.
Results
IEE can accurately and robustly quantify causality
To assess the numerical performance of measuring IntDC, we first applied the IEE algorithm on the following two-node Logistic dynamics
$\begin{cases}x_{t+1}=3.7\big{[}(1-\beta_{yx})x_{t}\left(1-x_{t}\right)+\beta_{yx}y_{t}\left(1-y_{t}\right)\big{]}+\varepsilon_{x,t},\\ y_{t+1}=3.7y_{t}\big{[}1-(1-\beta_{xy})y_{t}-\beta_{xy}x_{t}\big{]}+\varepsilon_{y,t},\end{cases}$ (8)
where $\varepsilon_{\cdot,t}$ represents independent Gaussian noises, and parameters $\beta_{xy}$ and $\beta_{yx}$ modulate the IntDC strength from $x$ to $y$ and from $y$ to $x$, respectively (see further details in SM).
IEE can accurately capture the numerical behavior of dynamical causality in the two-node Logistic system, especially IntDC rather than ObsDC. In Fig. 2(A), we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 00 to $0.3$. We observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots). In Fig. 2(B), we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 00 to $0.3$. $\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots). The gray dashed lines in Figs. 2(A-B) represent $0.01$ for reference. A detailed comparison with GC, TE and CCM can be found in figs. S1 and S2, which indicate that GC exhibited non-monotonic behavior, $\text{TE}[y\rightarrow x]$ slightly decreased with a false-negative issue, and $\text{CCM}[x\rightarrow y]$ had a false-positive problem. The mean values and standard deviations of the indices under 100 simulations are provided in tables S1-S4.
Within suitable parameter ranges, IEE demonstrates stability and robustness. In Figs. 2(D-G), we validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise amplitudes $\sigma$. The causal strength should exhibit a descending trend with decreasing $\beta_{yx}$ from $0.15$ to $0.05$ through $0.125,0.1,0.075$. The experiments support that IEE maintains the correct order under different parameter values with relatively stable variations.
IEE can eliminate the effects of confounding variables
IEE remains applicable and accurate even in presence of confounding variables. We simulated a three-node Logistic dynamics $(x_{t},y_{t},z_{t})$, where $z$ acted as a confounder (see details in SM). There were consistently non-zero causal effects from $z$ to $x$ and from $z$ to $y$, while $\beta_{xy}$ adjusted the causal strength from $x$ to $y$ (Fig. 2(C)). When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ was almost zero and significantly different from $\text{IEE}[z\rightarrow x]$ and $\text{IEE}[z\rightarrow y]$ (Fig. 2(C)); but conventional ObsDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig. S3). That is because an intervention on $x$, i.e. $\delta x$, will not induce $\delta y$ through $z$ under the IntDC framework. IEE, specially designed for measuring IntDC, remained unaffected by the confounder $z$ and yielded accurate results.
Figure 2: Performance of IEE on the Logistic systems. (A) and (B) used the two-node Logistic system. The $\beta_{xy}$ and $\beta_{yx}$ control the coupling coefficients from $x$ to $y$ and from $y$ to $x$, respectively. (A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 00 to $0.3$ over 100 simulations. $\text{IEE}[y\rightarrow x]$ (the red line with squares) increases monotonically, while $\text{IEE}[x\rightarrow y]$ (the blue line with dots) stays around zero. (B) is under $\beta_{xy}=0.1$, in which $\text{IEE}[x\rightarrow y]$ (the blue line with dots) is significantly positive. The gray dashed lines in (A) and (B) represent the constant $0.01$ as reference. (C) shows the performance of IEE in the three-node Logistic system, where $z$ acts as a confounding variable. When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ is almost zero correctly. IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives. The results for $\beta_{xy}=0.5$ is shown for a comparison. The gray dashed line is $0.05$ for reference. (D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise amplitudes $\sigma$ in the two-node logistic system, respectively. The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of $0.15$ (red), $0.125$ (blue), $0.1$ (black), $0.075$ (green), and $0.05$ (magenta). Lines denote mean values under 50 simulations, and shaded areas represent the standard deviation. IEE maintains correct orders of IntDC under different situations with relatively stable variations.
IEE can reconstruct causal networks with quantified causal strength
IEE has the capability to reconstruct the causal network structure and rank the causal influence between nodes in complex networks at the interventional level. We used a 10-node coupled Henon maps as an example, where each node serves as the dynamical cause to its subsequent node (Fig. 3(A)). Further details regarding the dynamics can be found in SM.
To verify the effectiveness of IEE in reconstructing causal networks, we calculated the Area Under Curve (AUC) values based on multiple simulated time series. The AUC value of IEE was $0.871\pm 0.008$, significantly higher than conventional ObsDC indices, such as GC with $0.707\pm 0.069$, TE with $0.816\pm 0.018$, and CCM with $0.837\pm 0.012$. The boxplots for AUC values are displayed in Fig. 3(B). AUCs for IEE exhibited a higher mean and a smaller variation compared to the others. Additionally, fig. S4 presents receiver operating characteristic (ROC) curves.
To show the capability for ranking the importance of causal influences, we conducted two tests. We first calculated the IEE from Node 1 to the other nine nodes. Consistently with the actual scenario, IEE displayed sequential decrease in IntDC originating from Node 1 across Nodes 2 to 10 (Fig. 3(C)). Then, we measured the IEE received by Node 7 from the other nine nodes. As expected, Nodes 1-6 exhibited ascending causal strengths on Node 7, whereas Nodes 8-10 had little influence on Node 7 (Fig. 3(D)). IEE accurately discerned the IntDCs, with Node 6 showing the strongest value. A comparison between IEE for the IntDC and the results from GC, TE, and CCM for the ObsDC can be found in figs. S5 and S6. GC and TE failed to rank the influence from Node 1 accurately (fig. S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 (fig. S6).
Figure 3: Performance of IEE on the 10-node coupled Henon-map dynamics. (A) illustrates the network structure of the dynamics. There are directed causality from each node to its subsequent node. (B) presents a boxplot of the Area Under Curve (AUC) values for the network reconstruction. IEE outperforms GC, TE, and CCM significantly in accuracy and stability. (C) demonstrates the IEE from Node 1 to the other nine nodes, while (D) shows the IEE received by Node 7 from the other nodes. IEE accurately captures the interventional flow and IntDC in cascade topology.
IEE can quantify IntDC effectively without requiring interventions or perturbations
IEE is specifically designed for calculating IntDC solely from observed time-series data without requiring additional perturbations to the system. To validate its accuracy, we conducted a comparison against true perturbed deviations using chaotic neural networks (chNNs) as a model system (?, ?).
The chNN comprises an output variable $\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network. We chose $N_{\text{node}}=10$ in simulations. Details on the dynamics can be found in SM. We denoted the observed time-series data from the stationary system without intervention as $\bm{x}_{\text{obs}}$. When the neuron $i$ was removed (set as a constant zero), the perturbed data were denoted as $\bm{x}_{\text{per}}^{(i)}$. Removing one node was taken as an intervention to the system. The stationary probability density functions of the $j$th neuron before and after the intervention were represented by $p_{j}(\bm{x}_{\text{obs}})$ and $p_{j}(\bm{x}_{\text{per}}^{(i)})$, respectively. We used the Kullback-Leibler divergence (KLD), i.e.
$D_{ij}$ $\triangleq\text{KLD}[p_{j}(\bm{x}_{\text{obs}})||p_{j}(\bm{x}_{\text{per}}^{(i)})]$
$=\iint p_{j}(\bm{x}_{\text{obs}})\log\frac{p_{j}(\bm{x}_{\text{obs}})}{p_{j}(\bm{x}_{\text{per}}^{(i)})}\mathop{}\!\mathrm{d}\bm{x}_{\text{obs}}\mathop{}\!\mathrm{d}\bm{x}_{\text{per}}^{(i)},$ (9)
to quantify the true influence of the intervention, i.e. IntDC, from $i$ to $j$. By conducting $100$ randomly simulated chNNs with interventions (10 removed nodes for each chNN), we recorded $1000$ KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely). In Fig. 4(A), gray dots are the $1000$ samples of causal edges between neurons, while the red line stands for the linear regression. We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$). Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) are shown in Figs. 4(B-D), which imply that GC and CCM cannot linearly reflect KLD. Furthermore, the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on $100$ chNNs are displayed in Fig. 4(E). IEE ($0.904\pm 0.024$) had the highest mean similarity with KLD, compared to GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). These results indicate that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
Figure 4: The performance of IEE when measuring the influence of perturbations on the chaotic neural networks (chNNs). The Kullback-Leibler divergence (KLD) is utilized to measure the true influence or IntDC between neurons. In (A-D) , $1000$ samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively. (E) displays the violin plots of cosine similarity between KLD and IEE/GC/TE/CCM. The mean values and standard variations of the cosine similarity are IEE ($0.904\pm 0.024$), GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). IEE can linearly reflect the KLD with high similarity, indicating its effectiveness in accurately quantifying IntDC without additional perturbed data on chNNs.
Application of IEE in inferring neural connectomes of C. elegans
We applied the IEE criterion to infer the neural connectomes of Caenorhabditis elegans (C. elegans), a model organism known for its comprehensively studied nervous system  (?). Calcium fluorescence imaging time series data from $31$ neurons with specific functions of a freely moving C. elegans were collected from Kato et al.  (?) (see Figs. 5(A-B) and fig. S7) . The Fig. 5(C) shows the cosine similarity matrix between different neurons. We clustered the neurons into $7$ clusters after data preprocessing (see details in Supplementary Text and table S5). Figure 5(D) presents the ground truth of neural connectomes in a directed graph, which was detected by electron microscopy in references  (?, ?, ?). We calculated the IEE between different clusters to infer the IntDC network. The ROC curve of IEE is shown in Fig. 5(F) with an AUC value $0.882$. The maximum Youden index, maximum concordance probability, and minimum distance to the point $(0,1)$ gave the same optimal operating point (OOP). At the OOP, Fig. 5(E) displays the inferred connectomic network by IEE. Comparisons with other ObsDC indices, i.e. GC/TE/CCM, revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs. S8-S10, and table S6). These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C. elegans.
Figure 5: Application of the IEE on inferring the neural connectomes of C. elegans. (A) is an illustration of C. elegans. (B) displays the calcium imaging time series of $31$ key neurons, with their names listed on the left and colors representing different clusters. (C) is the cosine similarity matrix between neurons, aiding the clustering process. (D) is the true connectomes between $7$ clusters, while (E) is the inferred causal network at the the optimal operating point (OOP) determined by IEE. The red edge represents a false positive, and the black dashed edges represent false negatives. (F) shows the ROC curve (the blue line) for IEE. The same OOP is obtained under three criteria, i.e. the maximum Youden index (the blue circle), the maximum concordance probability (the green dot), and the minimum distance to the point $(0,1)$ (the red cross). The shaded area around the ROC curve represents the $95\%$ confidence interval obtained through bootstrapping. The AUC value of IEE is $0.882$.
Application of IEE to COVID-19 transmission in Japan
We validated that IEE was a promising indicator for assessing the transmission dynamics of infectious diseases. Our study collected the daily confirmed COVID-19 cases from all $47$ prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig. S11).
First, we calculated the IEE indices for IntDC among different prefectures. IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa, Chiba, Saitama, Aichi, and Osaka. Remarkably, this ranking coincides with the geographical proximity and socio-economic connections to Tokyo. Kanagawa, Chiba and Saitama are in the same metropolitan area with Tokyo, while Aichi and Osaka are far from Tokyo but connected to Tokyo by expressway and Shinkansen.
Then, IEE was validated to coincide with the effective distance matrix $D^{\text{COVID}}$ between prefectures. The $D^{\text{COVID}}\in\mathbb{R}^{47\times 47}$ was non-symmetric and designed by incorporating factors such as geometric distances, human mobility, population sizes, and infectious rates across prefectures (based on the gravity model  (?, ?), see Supplementary Text and fig. S11 for details). The element $D_{ij}^{\text{COVID}}$ served as a benchmark for quantifying COVID-19 transmission from prefecture $i$ to $j$. The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (Figs. 6(A-B)). This correlation remained significant for other regions such as Aichi, Hokkaido, Fukuoka, and Okinawa (see fig. S12).
Further, IEE outperformed ObsDC indices in practicality. Notably, the absolute value of the Pearson correlation coefficient (PCC) between IEE and $\ln D^{\text{COVID}}$ was $0.724$ substantially higher than those obtained by GC $0.161$, TE $0.284$, and CCM $0.568$ (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig. 6(C)). Table 3 presented more detailed comparisons to support the effectiveness of IEE (with Tokyo for specific illustration in fig. S13).
These results underscore the reliability and efficacy of IEE in providing quantitative insights into COVID-19 transmission dynamics in Japan. Importantly, IEE performs as a simple and valuable tool for causal analyses solely from daily confirmed time-series data, without the need for complex models considering various factors such as geodesic distances, human mobility, population sizes, and infectious rates.
Figure 6: Application of the IEE on the COVID-19 transmission in Japan and the gene regulatory network (GRN) of circadian rhythms. (A) shows that IEE values from Tokyo to other prefectures are highly linear-correlated with $\ln D^{\text{COVID}}$, where $D^{\text{COVID}}$ is the effective distance obtained from the gravity model in consideration of various factors. (B) is the result for IEE values from Osaka to the other prefectures. (C) is the boxplot for $|\text{PCC}(c_{ij}^{\text{Inf}},\ln D_{ij}^{\text{COVID}})|$ with average over fixed $i$, where PCC is the Pearson correlation coefficient and $c^{\text{Inf}}$ is the inferred causal strengths by IEE, GC, TE, or CCM. The PCC linearity of IEE ($0.724\pm 0.116$) is significantly higher than that of GC ($0.161\pm 0.099$), TE($0.284\pm 0.168$), and CCM ($0.568\pm 0.174$). (D) shows the GRN surrounding Clock, a key circadian gene. (E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ObsDC indices GC/TE/CCM.
Application of IEE to investigating circadian rhythms
We investigated the gene regulatory networks (GRNs) related to key genes on the circadian rhythm. The time series of gene expressions were measured by microarray from cultured rat cells  (?, ?). Through decades of molecular and genetic studies, many key circadian genes, such as Clock, Bmal1(Arntl), Dec1(Bhlhb2), Dec2(Bhlhb3), Cry1, Cry2, Per1, Per2, Per3, have been identified on mammals  (?, ?). Figure 6(D) displayed the GRN around Clock at the protein level, where the transcription factor Clock is phosphorylated by PFK family genes. The ROC curves and AUC values for IEE/GC/TE/CCM were showed in Fig. 6(E). IEE designed for quantifying IntDC presented a higher AUC value ($0.737$) than ObsDC indices ($0.490$ for GC, $0.610$ for TE, and $0.644$ for CCM). In fig. S14, we also tested the indices on the GRN surrounding Cry1/Cry2, where IEE outperformed the others.
Concluding remarks and perspective
In summary, we have established a framework of IntDC based on dynamical systems theory and introduced the IEE criterion to quantify IntDC in this study. IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?). Moreover, IEE is able to infer IntDC solely from observational time-series data, without requiring additional perturbations to the system. This property makes IEE particularly suitable for non-intervention systems. Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the high effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks. Furthermore, IEE serves as an effective tool for evaluating and ranking the causal dependence between variables within a dynamical system. Through real-world examples, we illustrated the promising applicability of IEE in diverse fields such as regulatory inference and disease transmission studies.
There are several compelling issues that warrant further investigation and future development.
(1) The theory of IntDC and the IEE criterion can be extended to analyze causal relations between groups of variables. Specifically, in Eq. (6), the variables $x$ and $y$ can be generalized to high-dimensional vectors, enabling the investigation of complex interactions among multiple variables.
(2) The data in computing IEE can be non-uniformly sampled in the temporal space. We used the time-delayed embedding Eq. (2) and Eq. (3) in this paper, while according to the general Takens’ embedding theorem, any $L+1$ observations of $x_{t}$ and $y_{t}$ can form a topological diffeomorphism to $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively. Equation (6) shows that only neighbors in the embedding space matter. Thus, IEE algorithm can be extended to infer causality using diverse datasets sampled at different intervals.
(3) To distinguish direct and indirect causality, we can generalize IEE to its conditional version
$\text{cIEE}[x\rightarrow y|z]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (10)
where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delay vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$. Equation (10) measures the direct IntDC from $x$ to $y$ conditioned on $z$. Further, together with the well-known PC algorithm  (?), cIEE is one feasible way to remove high-order indirect causal edges iteratively and get the direct causal network.
(4) Moreover, causal inference with deep learning technology represents a frontier in research. Notable works include CausalEGM  (?) and Causalformer (?). Combining IntDC and neural networks is a promising field. Causality theory brings interpretability for artificial intelligence, while big model can provide new discoveries in causal relationships. We are now entering a brilliant “big causality era”.
Acknowledgments
Funding:
This work was supported by the National Key R&D Program of China 2022YFC2704604, the National Natural Science Foundation of China No. 12301620 and No. 32100509, the Moonshot R&D Grant Number JPMJMS2021, Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), JSPS KAKENHI Grant Number JP20H05921, the Cross-ministerial Strategic Innovation Promotion Program (SIP), and the 3rd period of SIP “Smart energy management system” Grant Number JPJ012207.
Author contributions:
Conceptualization: J.S., K.A., L.C., and W.L.
Methodology: J.S., K.A., L.C., and W.L.
Investigation: J.S., Y.L., J.Z., and S.L.
Visualization: J.S., Y.L., J.Z., S.L.
Supervision: K.A., L.C., and W.L.
Writing – original draft: J.S, Y.L., J.Z., and L.C.
Writing – review & editing: All authors.
Competing interests:
The authors declare no competing interest.
Data and materials availability:
The code for IEE is available at https://github.com/smsxiaomayi/IEE.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S14
Tables S1 to S6